OpenStudy (anonymous):

{Could someone help with this?} Point A(–4, 2) is reflected over the line x = 3 to create the point A'. What are the coordinates of A'? A. (–4, –1) B. (–4, 8) C. (–1, 2) D. (10, 2)

OpenStudy (anonymous):

 now all you have to do in reflection is plot the point on the other side of the line x=3, at the same distance as it was... since the distance on the one side was 7 units.. you just add 2*7 =14 to the x coordinate and the point will automatically go as intended... ur new point will be (2*7-4,2) = (14-4,2) = (10,2)